OpenStudy (anonymous):

How would I find the cumulative distribution function (CDF) of the following: f(x) = e^(-x^2/2a^2) * (x/a^2) ?

hartnn (hartnn):

assuming f(x) as the density function, CDF is : \(\Large CDF = \int \limits_{-\infty }^x f(x)dx\)

hartnn (hartnn):

to solve that integral, please try substitution method u = -x^2/2a^2 and let me know if you get stuck at any step :)

OpenStudy (anonymous):

Thanks so much hartnn, that really helps. I'll hit you up if I hit another snag in this statistics homework I'm doing!

hartnn (hartnn):

sure, welcome ^_^